What is the present value of $100 ten years from now if the simple rate of interest is0.26? Round your final answer to two decimal places. Question 2 2 pts A deposit of $19 is made at the beginning of each year for 14 years, what is the future value of al deposits if simple rate of interest is 0.09? Round your final answer to two decimal places.
Answer #1

a. Using the Present Value formula as:
PV = FV 1/(1+R)n
PV = 100/(1+0.0026)10

= $97.44


b. To find the Future value of Equal Investments of $19 made at the beginning of each year for 14 years at 0.09 Simple Interest rate using formula:
FV = PV (1xR)n
FV = 19 (1x0.09)14
= $267.8